A 47.3 kg wheelchair rolls down a 17.4 m long ramp with an angle of inclination of 20.3° in 7.1 s. What is the power of gravity on the wheelchair?



1
Expert's answer
2021-06-28T16:32:03-0400

Solution.

"m=47.3kg;"

"S=17.4m;"

"\alpha=20.3^o;"

"t=7.1s;"

"S=\dfrac{at^2}{2}\implies a=\dfrac{2S}{t^2};"

"a=\dfrac{2\sdot17.4}{7.1^2}=0.69m/s^2;"

"a_f=\sqrt{a^2+g^2+2gacos\alpha}=\sqrt{0.69^2+9.8^2+2\sdot0.69\sdot9.8\sdot0.9391}=10.45m/s^2;"

"F=ma_f;"

"F=47.3\sdot 10.45=494.3N;"

Answer: "F=494.3N."